The electric field is measured to be 22.7 N/C due to a point charge of 6.1 nC. How far away was the electric field measured?
1
Expert's answer
2020-02-27T09:45:33-0500



The field of point charge is described by the formula

"E=k \cdot \frac{|q|}{r^2}"

Where will we write

"r=\sqrt{\frac{k \cdot |q|}{E}}=\sqrt{\frac{9 \cdot 10^9 \cdot 6.1 \cdot 10^{-9}}{22.7}}=1.555 m"